OpenStudy (waheguru):

The lenght from the center of teh circle to a point on the circle is r = square root of 40 what is the length of the radius the answer is 2 and squareroot 10?

OpenStudy (anonymous):

\[ \sqrt{40} = \sqrt{4}\times\sqrt{10}=2\sqrt{10} \]

OpenStudy (waheguru):

oh THANK YOU

OpenStudy (waheguru):

:)

OpenStudy (waheguru):

but the first number has to be a perfec square right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

You can pull out any perfect squares.

OpenStudy (anonymous):

Since \(40=2\cdot 2\cdot 2\cdot 5\) you can pull out the \(2\cdot 2\), and he \(2\cdot 5\) remains inside.